OpenStudy (anonymous):

A line passes through the point (-3,5) and has a slope of -5 . Write an equation for this line.

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] put \[x_1=-3, y_1=5, m=-5\]

OpenStudy (anonymous):

so whats the equation

OpenStudy (anonymous):

@satellite73